OpenStudy (anonymous):

What is the equation of a line passing through (-3, 7) and having a slope of -1/5 ?

OpenStudy (vishweshshrimali5):

The equation of a line passing through \((x_1,y_2)\) and having a slope m, is, \[\large{(y-y_1) = m(x-x_1)}\]

OpenStudy (anonymous):

y = -5x + 22 y = 5x + 22 y=-1/5x+32/5 y=1/5x+32/5 y=-1/5x+22 which is right?

OpenStudy (anonymous):

;-;?

OpenStudy (vishweshshrimali5):

I have given you the general equation. In that equation, put m = -1/5 and (x1,y1) = (-3,7)

OpenStudy (anonymous):

idk how to do this ive never learned it and i have 27 minutes to pass this semester test ;-;

OpenStudy (anonymous):

im not just trying to get answers id love to learn it but i dont have time right now i have to pass so i can go to the military

OpenStudy (vishweshshrimali5):

Okay after putting the values and solving you will get: C is correct.

OpenStudy (anonymous):

thank you like alot..

OpenStudy (vishweshshrimali5):

No problem and best of luck! :)